OpenStudy (anonymous):

The mass of an object: a.is 0.1 kg if it weighs 98 N at g=9.8 m/s^2.. b.is 10kg if it weighs 98 N at g=9.8 m/s^2. c.of an object is 1kg if it weighs 98 N at g=9.8 m/s^2. d.is 100 kg if it weighs 98N at g=9.8 m/s^2.

OpenStudy (aaronq):

\(F=ma\) \(F=98N=m(9.8/ms^2)\) \(1~N=1~kg*m/s^2\) \(m=\dfrac{98~N}{9.8~m/s^2}=\dfrac{98~kg*m/s^2}{9.8~m/s^2}=10~kg\)

OpenStudy (calculusfunctions):

As @aaronq said: Weight = mass X acceleration due to gravity In other words, W = mg